# The thermodynamic structure and large-scale structure filament in MACS J0717.5+3745

### v2

@@
11institutetext: Department of Theoretical Physics and Astrophysics, Masaryk University. Kotlářská 2, Brno, 611 37, Czech Republic 22institutetext: Department of Physics, Graduate School of Advanced Science and Engineering, Hiroshima University Kagamiyama, 1-3-1 Higashi-Hiroshima, 739-8526, Japan 33institutetext: NASA Goddard Space Flight Center, Code 662, Greenbelt, MD 20771, USA 44institutetext: Department of Astronomy, University of Maryland, College Park, MD 20742-2421, USA 55institutetext: ESA/ESTEC, Keplerlaan 1, 2201 AZ Noordwijk, The Netherlands 66institutetext: Academia Sinica Institute of Astronomy and Astrophysics (ASIAA), No. 1, Section 4, Roosevelt Road, Taipei 106216, Taiwan 77institutetext: SRON Netherlands Institute for Space Research, Niels Bohrweg 4, 2333CA Leiden, The Netherlands 88institutetext: Leiden Observatory, Leiden University, PO Box 9513, 2300 RA Leiden, The Netherlands 99institutetext: Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo, Kashiwa, Chiba 277-8583, Japan 1010institutetext: University of Pennsylvania, 209 S. 33rd St., Philadelphia, PA 19014, USA $1$institutetext: Kapteyn Astronomical Institute, University of Groningen, Landleven 12, 9747 AD, Groningen, The Netherlands 1212institutetext: Laboratoire Lagrange, Université Côte d’Azur, Observatoire de la Côte d’Azur, CNRS, Blvd de l’Observatoire, CS 34229, 06304 Nice cedex 4, France 1313institutetext: Department of Physics, McGill University, 3600 University Street Montreal, QC, H3A 2T8, Canada 1414institutetext: National Radio Astronomy Observatory, 520 Edgemont Rd., Charlottesville VA 22903, USA 1515institutetext: European Southern Observatory, Karl-Schwarzschild-Straße 2, 85748 Garching bei München, Germany 1616institutetext: Center for Astrophysics, Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA 1717institutetext: Department of Astronomy, University of Virginia, P.O. Box 400325, Charlottesville, VA 22904, USA
@@
J. P. Breuer E-mail: jeanpaul.breuer@gmail.com1122    N. Werner 11    T. Plšek 11    F. Mernier 334455    K. Umetsu 66    A. Simionescu 778899    M. Devlin 1010    L. Di Mascolo 11111212    T. Dibblee-Barkman 1313    S. Dicker 1010    B. S. Mason 1414    T. Mroczkowski 1515    C. Romero 1616    C. L. Sarazin 1717    J. Sievers 1313
@@
We present the results of Chandra and XMM-Newton X-ray imaging and spatially-resolved spectroscopy, as well as new MUSTANG2 90 GHz observations of the thermal Sunyaev-Zeldovich effect from MACS J0717.5+3745, an intermediate redshift ($z=0.5458$) and exceptionally massive ($3.5\pm 0.6\times 10^{15}$ M) Frontier Fields cluster experiencing multiple mergers and hosting an apparent X-ray bright large scale structure filament. Thermodynamical maps are produced from Chandra, XMM-Newton, and ROSAT data using a new method for modeling the astrophysical and instrumental backgrounds. The temperature peak of $24\pm 4$ keV is also the pressure peak of the cluster and closely correlates spatially with the Sunyaev-Zeldovich peak from the MUSTANG2 data. The cluster center hosts shock fronts to the north and south, for which we report estimates for the shock Mach numbers of $\mathcal{M}=1.6\pm 0.4$ and $\mathcal{M}=1.9\pm 0.3$, respectively. Bayesian X-ray Analysis methods were used to disentangle different projected spectral signatures for the filament structure, with Akaike and Bayes criteria being used to select the most appropriate model to describe the various temperature components. We report an X-ray filament temperature of $3.1_{-0.3}^{+0.6}$ keV and a density $(3.78\pm 0.05)\times 10^{-4}\,{\rm cm^{-3}}$, corresponding to an overdensity of $\sim 400$ relative to the critical density of the Universe. We estimate the hot gas mass of the filament to be $\sim 6.1\times 10^{12}\leavevmode\nobreak\ \rm M_{\odot}$, while its total projected weak lensing measured mass is $\sim(6.8\pm 2.7)\times 10^{13}\leavevmode\nobreak\ \rm M_{\odot}$, indicating a hot baryon fraction of 4–10%.
@@ 1 Introduction
Current cosmological models and numerical simulations predict that the majority of the missing baryons in our Universe sit in the faint galaxy cluster outskirts and the interconnecting filaments of the cosmic web (Cen & Ostriker, 1999; Davé et al., 2001). This warm-hot intergalactic medium (WHIM) contributes to the growth of galaxy clusters through a slow constant accretion, as the in-falling filamentary gas virializes within the gravitational potential wells of massive clusters.
Only a small number of X-ray bridges and filaments have been studied so far. This is partially because the diffuse WHIM remains a complicated challenge to detect with today’s instruments. Numerical cosmological simulations predict WHIM temperatures of $10^{5}$ to $10^{7}$ K and densities of 10-7 to 10-4 cm-3 (Cen & Ostriker, 2006; Haider et al., 2016); in other words, features in the WHIM are inherently characterized by soft X-ray emission, low surface brightness, and poor signal to noise ratios due to dominant contributions to the signal from both astrophysical and instrumental backgrounds. Needless to say, much care needs to be taken when modelling the spectra of these complex regions to account for all of the different contributions from other emission signatures.
Current observational information about filaments is derived from either absorption or emission. Observations in emission have so far been limited to a small number of systems, MACS J0717.5+3745 (Ebeling et al., 2004), Abell 399/401 (Sakelliou & Ponman, 2004), Abell 222/223 (Werner et al., 2008), Abell 2811 (+offset)/2804/2801 (Sato et al., 2010), Abell 3558/3556 (Mitsuishi et al., 2012), Abell 2744 (Eckert et al., 2015a), Abell 3391/3395 (Alvarez et al., 2018; Reiprich et al., 2021; Veronica et al., 2022, 2024), Abell 2029/2033 (Mirakhor et al., 2022), Abell 98N/98S (Sarkar et al., 2022), and Abell 3667/3651 (Dietl et al., 2024). Temperature measurements were obtained only for four of these systems. For two of the four, it has been reported that the filament emission is dominated by the emission of the Intracluster Medium (ICM), as evidenced by the excessively large temperatures.
@@
One of the most extensively studied merging galaxy clusters in almost every available wavelength is MACS J0717.5+3745 (RA 07h17m32.1, DEC +3745’21”), an extremely massive, intermediate redshift ($z=0.5458$), Hubble Frontier Fields Cluster (Ebeling et al., 2001). MACS J0717.5+3745 is also one of the most complex galaxy clusters known to date, having hosted at least four different sub-cluster collisions. The collective mass from the many ingested dark matter halos makes this a great target for gravitational lensing studies and additional exploration of local substructures via the Sunyaev-Zeldovich effect due to the high temperatures. The complex mergers have also resulted in a complex ICM morphology, characterized by perturbed plasma discontinuities, stripped atmospheres, shock fronts, and cold fronts, and results in many peculiar features in the X-ray morphology and radio surface brightness features. Besides the gravitational lensing study, Jauzac et al. (2018), also looked at many of these X-ray substructures in detail. Most intriguing is the prominent X-ray bridge located in the S-SE of the cluster, which has previously been studied by van Weeren et al. (2016, 2017) using Chandra data.
@@ 2.1 Chandra Observations
The data from four Chandra pointings (see Tab. 1) were reprocessed using repro from the level-1 event lists with the standard software packages using the most recent versions of CIAO and CALDB (versions 4.16 and 4.11.0). The good time intervals (GTI) of the observations after filtering periods of flaring are summarized in Tab. 1. All observations were used for creating the surface brightness images; however, only observations 4200, 16235, and 16305 were considered in further spectral analysis, as observation 1655 experienced residual soft proton flaring contamination which would bias spectral analysis. Totalling all available Chandra observations gives approximately 228.9 ks of observing time; however, with the removal of observation 1655, only 211.8 ks were used for creating the data products. Because of the low signal-to-noise ratio, blank-sky background files were explicitly not used in favour of using epoch-specific stowed background files normalized by the data over background count ratio in the 10-12 keV energy range. When dealing with extremely faint X-ray substructures, the stowed background files give additional control over the systematics in the background. Merged broadband images were produced in the 0.5 to 7 keV band, shown in Fig. 1 (left). Unless otherwise specified, the spectra of all regions were extracted from each Chandra observation using the specextract routine from the latest version of CIAO (version 4.16.0).
@@ 2.2 XMM-Newton Observations
Complementary to the Chandra data, we also retrieved archival XMM-Newton observations of MACS J0717.5+3745, as listed in Tab. 1. Three pointings showing sufficient data quality were downloaded and reduced using both the XMM Science Analysis System (SAS) software (v20.0.0) and the current calibration files (CCF) by following the procedure detailed in the Extended Source Analysis Software (ESAS v20.0.0) cookbook. To get cleaned event files for each detector, we used the standard emchain and mos-filter routines on the MOS data, and the epchain and pn-filter routines on the pn data, which removes periods of anomalously high count rate by calculating GTIs through the espfilt routine. The total GTI for each observation are also shown in Tab. 1.
@@
Figure 1: Images of the full FOV of Chandra (left) and XMM-Newton (right) of MACS J0717.5+3745. The surface brightness excess to the west/south-west are several foreground structures in between us and the cluster. The green dashed circle indicates the approximate $R_{200}$ of the cluster. Intensity scale is in flux per pixel (0.492$\times$0.492 arcsec for Chandra, and 4.1$\times$4.1 arcsec for XMM pn).
Figure 2: The Chandra (left) and XMM-Newton (right) images of MACS J0717.5+3745, restricted to the same field of view as the Sunyaev-Zeldovich data and thermodynamical maps, visible as the box in Fig. 1. Intensity scale is in flux per pixel (0.492$\times$0.492 arcsec for Chandra, and 4.1$\times$4.1 arcsec for XMM pn). Part of the XMM-Newton image is over-saturated so that the filament structure is more visible. The white lines indicate the regions used for modelling the spectra of the outskirts of the cluster and the group together with the filament. The green dashed circle indicates the approximate $R_{200}$ of the cluster.
Table 1: Summary of the data, excluding overly contaminated observations. GTI is the net exposure good time intervals for all instruments after cleaning the data. ()(*)( ∗ ) Not used.
@@ 2.4 MUSTANG2 Sunyaev-Zeldovich Observations
MACS J0717.5+3745 was observed by the 100-meter Green Bank Telescope (GBT) under project codes AGBT17A_340 and AGBT17B_266 (in semesters 2017A & 2017B, respectively) using the MUSTANG-2 instrument, a 215-element continuum bolometer array operating at 90 GHz (Dicker et al., 2014). The observations were conducted using Lissajous daisy scans with radii of 3′ and ${\prime}$5 (Romero et al., 2020, 2023a). The data processing was identical to that in, for example, Romero et al. (2020). The resulting map, produced using the MUSTANG-2 pipeline, is shown in the lower left panel of Figure 9. The total integration time on target was $2.6$ hours and the resulting map has a resolution of $10^{\prime\prime}$, with an RMS noise level of $\sim 28\,{\rm\mu Jy/bm}$. Besides the astrophysical contaminants and the instrumental noise, MUSTANG2 data is additionally affected by large scale filtering of the astrophysical signal, which affects the imaging quality.
@@ 3.1 Chandra Stowed Instrument Background
Of the available background datasets, the total stowed background merged exposure time was 2.91 Ms from the year 2000, 2.57 Ms from the year 2005, and 1.68 Ms for the year 2009. Spectra were extracted from the full field of view of each available year and were modelled following a similar process as described in Siemiginowska et al. (2010) for Chandra and Mernier et al. (2015) for XMM-Newton, with independent fits for each of the detectors. Special care was also taken regarding the Faint or Very Faint imaging modes, resulting in modified models that involve these different cases. The stowed background data was modelled in the 0.7 keV to 10 keV band for both the ACIS-I and ACIS-S detectors. The continuum emission is defined by a broken power-law for the Faint imaging submode and by a power-law in the Very Faint imaging submode, while all instrumental lines were modelled using Gaussians and frozen to their best fits. Bartalucci et al. (2014) noted that the shape of the particle background is fairly consistent, and only the normalizations in different epochs are changing from year to year; therefore, we refrain from including redundant information for all years, and give detailed results of the continuum emission and instrumental lines in Tab. 2. Elements corresponding to line energies used in our background model were documented in other work (Bearden & Burr, 1967; Suzuki et al., 2021). The line energies which did not exactly correspond to reported lines were marked in the table with ‘*’, while the 2.6 keV line was marked with ‘**∗ ∗’ as it was not previously mentioned in other reference material. This line in particular showed up only in the ACIS-I stowed backgrounds and is unlikely to be an instrumental line, rather some artifact of the response itself. It was included in the ACIS-I model as an empirical improvement for completeness. All Chandra spectral fits reported in this paper were performed in the 0.7 to 7.0 keV range.
@@
[…]
2.6∗∗ - - -
[…]

### v3

@@
11institutetext: Department of Theoretical Physics and Astrophysics, Masaryk University. Kotlářská 2, Brno, 611 37, Czech Republic 22institutetext: Department of Physics, Graduate School of Advanced Science and Engineering, Hiroshima University Kagamiyama, 1-3-1 Higashi-Hiroshima, 739-8526, Japan 33institutetext: NASA Goddard Space Flight Center, Code 662, Greenbelt, MD 20771, USA 44institutetext: Department of Astronomy, University of Maryland, College Park, MD 20742-2421, USA 55institutetext: ESA/ESTEC, Keplerlaan 1, 2201 AZ Noordwijk, The Netherlands 66institutetext: Academia Sinica Institute of Astronomy and Astrophysics (ASIAA), No. 1, Section 4, Roosevelt Road, Taipei 106216, Taiwan 77institutetext: SRON Netherlands Institute for Space Research, Niels Bohrweg 4, 2333CA Leiden, The Netherlands 88institutetext: Leiden Observatory, Leiden University, PO Box 9513, 2300 RA Leiden, The Netherlands 99institutetext: Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo, Kashiwa, Chiba 277-8583, Japan 1010institutetext: University of Pennsylvania, 209 S. 33rd St., Philadelphia, PA 19014, USA 1111institutetext: Kapteyn Astronomical Institute, University of Groningen, Landleven 12, 9747 AD, Groningen, The Netherlands 1212institutetext: Laboratoire Lagrange, Université Côte d’Azur, Observatoire de la Côte d’Azur, CNRS, Blvd de l’Observatoire, CS 34229, 06304 Nice cedex 4, France 1313institutetext: Department of Physics, McGill University, 3600 University Street Montreal, QC, H3A 2T8, Canada 1414institutetext: National Radio Astronomy Observatory, 520 Edgemont Rd., Charlottesville VA 22903, USA 1515institutetext: European Southern Observatory, Karl-Schwarzschild-Straße 2, 85748 Garching bei München, Germany 1616institutetext: Center for Astrophysics, Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA 1717institutetext: Department of Astronomy, University of Virginia, P.O. Box 400325, Charlottesville, VA 22904, USA
@@
J. P. Breuer E-mail: jeanpaul.breuer@gmail.com    N. Werner    T. Plšek    F. Mernier    K. Umetsu    A. Simionescu    M. Devlin    L. Di Mascolo    T. Dibblee-Barkman    S. Dicker    B. S. Mason    T. Mroczkowski    C. Romero    C. L. Sarazin    J. Sievers
@@
We present the results of Chandra and XMM-Newton X-ray imaging and spatially resolved spectroscopy, as well as new MUSTANG2 90 GHz observations of the thermal Sunyaev-Zeldovich effect from MACS J0717.5+3745, an intermediate redshift ($z=0.5458$) and exceptionally massive ($3.5\pm 0.6\times 10^{15}$ M) Frontier Fields cluster experiencing multiple mergers and hosting an apparent X-ray bright large-scale structure filament. Thermodynamical maps are produced from Chandra, XMM-Newton, and ROSAT data using a new method to model the astrophysical and instrumental backgrounds. The temperature peak of $24\pm 4$ keV is also the pressure peak of the cluster and is spatially closely correlated with the Sunyaev-Zeldovich peak from the MUSTANG2 data. We characterize a potential shock candidate at the cluster center, based on the sharp temperature and pressure gradient, and quantify its temperature-derived Mach number in various directions to span a range of $\mathcal{M}=(1.7-2.0)\pm 0.3$. Bayesian X-ray Analysis methods were used to disentangle different projected spectral signatures for the filament structure, with Akaike and Bayes criteria being used to select the most appropriate model to describe the various temperature components. We report an X-ray filament temperature of $3.1_{-0.3}^{+0.6}$ keV and a density $(3.78\pm 0.05)\times 10^{-4}\,{\rm cm^{-3}}$, corresponding to an overdensity of $\sim 400$ relative to the critical density of the Universe. We estimate the hot gas mass of the filament to be $\sim 6.1\times 10^{12}\penalty 10000\ \rm M_{\odot}$, while its total projected weak lensing measured mass is $\sim(6.8\pm 2.7)\times 10^{13}\penalty 10000\ \rm M_{\odot}$, indicating a hot baryon fraction of 4–10%.
@@ 1 Introduction
Current cosmological models and numerical simulations predict that the majority of the missing baryons in our Universe sit in the faint galaxy cluster outskirts and the interconnecting filaments of the cosmic web (Cen & Ostriker, 1999; Davé et al., 2001). This warm-hot intergalactic medium (WHIM) contributes to the growth of galaxy clusters through a slow, constant accretion, as the infalling filamentary gas virializes within their gravitational potential wells.
Only a small number of X-ray bridges and filaments have been studied so far. This is partially because the diffuse WHIM remains a complicated challenge to detect with today’s instruments. Numerical cosmological simulations predict WHIM temperatures of $10^{5}$ to $10^{7}$ K and densities of 10-7 to 10-4 cm-3 (Cen & Ostriker, 2006; Haider et al., 2016); in other words, features in the WHIM are inherently characterized by soft X-ray emission, low surface brightness, and poor signal to noise ratios due to dominant contributions to the signal from both astrophysical and instrumental backgrounds. Needless to say, much care needs to be taken when modelling the spectra of these complex regions to account for all of the different contributions from other emission components.
Current observational information about filaments is derived from either absorption or emission. Observations in emission have so far been limited to a small number of systems, MACS J0717.5+3745 (Ebeling et al., 2004), Abell 399/401 (Sakelliou & Ponman, 2004), Abell 222/223 (Werner et al., 2008), Abell 2811 (+offset)/2804/2801 (Sato et al., 2010), Abell 3558/3556 (Mitsuishi et al., 2012), Abell 2744 (Eckert et al., 2015a), Abell 3391/3395 (Alvarez et al., 2018; Reiprich et al., 2021; Veronica et al., 2022, 2024), Abell 2029/2033 (Mirakhor et al., 2022), Abell 98N/98S (Sarkar et al., 2022), Abell 3667/3651 (Dietl et al., 2024), and A3530/32 and A3528-N/S (Migkas et al., 2025). Temperature measurements were obtained only for five of these systems. For two of the five, it has been reported that the filament emission is dominated by the emission of the Intracluster Medium (ICM), as evidenced by the excessively large temperatures.
@@
One of the most extensively studied merging galaxy clusters in almost every available wavelength is MACS J0717.5+3745 (RA 07h17m32.1, DEC +3745’21”), an extremely massive, intermediate redshift ($z=0.5458$), Hubble Frontier Fields Cluster (Ebeling et al., 2001). MACS J0717.5+3745 is also one of the most complex galaxy clusters known to date, having hosted at least four different sub-cluster collisions. The collective mass from the many ingested dark matter halos makes this a great target for gravitational lensing studies and additional exploration of local substructures via the Sunyaev-Zeldovich effect due to the high temperatures. The mergers have also resulted in a complex ICM morphology, characterized by discontinuities, such as shock fronts and cold fronts, as well as many peculiar features in the X-ray and radio bands. Besides the gravitational lensing study, Jauzac et al. (2018), also looked at many of these X-ray substructures in detail. Most intriguing is the prominent X-ray bridge located in the S-SE of the cluster, which has previously been studied by van Weeren et al. (2016, 2017) using Chandra data.
@@ 2.1 Chandra Observations
The data from four Chandra pointings (see Table 1) were reprocessed using repro from the level-1 event lists with the standard software packages using the most recent versions of CIAO and CALDB (versions 4.16.0 and 4.11.0). The good time intervals (GTI) of the observations after filtering periods of flaring are summarized in Table 1. All observations were used for creating the surface brightness images; however, only observations 4200, 16235, and 16305 were considered in further spectral analysis, as observation 1655 experienced residual soft proton flaring contamination which would bias spectral analysis. Totalling all available Chandra observations gives approximately 228.9 ks of observing time; however, with the removal of observation 1655, only 211.8 ks were used for spectral analysis. Because of the low signal-to-noise ratio, blank-sky background files were explicitly not used in favour of using epoch-specific stowed background files normalized by the data over background count ratio in the 10-12 keV energy range. When dealing with extremely faint X-ray substructures, the stowed background files give additional control over the systematics in the background. Merged broadband images were produced in the 0.5 to 7 keV band, shown in Fig. 1 (left).
@@ 2.2 XMM-Newton Observations
Complementary to the Chandra data, we also retrieved archival XMM-Newton observations of MACS J0717.5+3745, as listed in Table 1. Three pointings showing sufficient data quality were downloaded and reduced using both the XMM Science Analysis System (SAS) software (v20.0.0) and the current calibration files (CCF) by following the procedure detailed in the Extended Source Analysis Software (ESAS v20.0.0) cookbook. To get cleaned event files for each detector, we used the standard emchain and mos-filter routines for the MOS data, and the epchain and pn-filter routines for the pn data. mos-filter and pn-filter remove periods of anomalously high count rate by calculating GTIs through the espfilt routine. The total GTI for each observation are also shown in Table 1.
@@
Figure 1: Images of the full FOV of Chandra (left) and XMM-Newton (right) of MACS J0717.5+3745. The surface brightness excess to the west/south-west are several foreground structures in between us and the cluster, indicated by the ‘Foreground Emission’ label. The XMM-Newton image is over-saturated in the cluster center so that this can be seen more clearly. The green dashed circle indicates the approximate $R_{200}$ of the cluster. Intensity scale is in flux per pixel ($\times0.492$ arcsec for Chandra, and $\times4.1$ arcsec for XMM pn).
Figure 2: The Chandra (left) and XMM-Newton (right) images of MACS J0717.5+3745, restricted to the same field of view as the Sunyaev-Zeldovich data and thermodynamical maps, visible as the box in Fig. 1. Intensity scale is in flux per pixel ($\times0.492$ arcsec for Chandra, and $\times4.1$ arcsec for XMM pn). Part of the XMM-Newton image is over-saturated so that the filament structure is more visible. The white lines indicate the regions used for modelling the spectra of the outskirts of the cluster and the group together with the filament. The green dashed circle indicates the approximate $R_{200}$ of the cluster.
Table 1: Summary of the data, excluding overly contaminated observations. GTI is the net exposure good time intervals for all instruments after cleaning the data. ()(*) Not used.
@@ 2.4 MUSTANG2 Sunyaev-Zeldovich Observations
[…]
@@ 3.1 Chandra Stowed Instrument Background
Of the available background datasets, the total stowed background merged exposure time was 2.91 Ms for the year 2000, 2.57 Ms for 2005, and 1.68 Ms for 2009. Spectra were extracted from the full field of view of each available year and were modelled following a similar process as described in Siemiginowska et al. (2010) for Chandra and Mernier et al. (2015) for XMM-Newton, with independent fits for each of the detectors. Special care was also taken regarding the Faint or Very Faint imaging modes, resulting in modified models that involve these different cases. The stowed background data was modelled in the 0.7 keV to 10 keV band for both the ACIS-I and ACIS-S detectors. The continuum emission is defined by a broken power-law for the Faint imaging submode and by a power-law in the Very Faint imaging submode, while all instrumental lines were modelled using Gaussians and frozen to their best fits. Bartalucci et al. (2014) noted that the shape of the particle background is fairly consistent, and only the normalizations in different epochs are changing from year to year; therefore, we refrain from including redundant information for all years, and give detailed results of the continuum emission and instrumental lines in Table 2. Elements corresponding to line energies used in our background model were documented in other work (Bearden & Burr, 1967; Suzuki et al., 2021).
The line energies which did not exactly correspond to reported lines were marked in the table with ‘*’, while the 2.6 keV line was marked with ‘**’ as it was not previously mentioned in other reference material. This line in particular showed up only in the ACIS-I stowed backgrounds and is unlikely to be an instrumental line, rather some artifact of the response itself. It was included in the ACIS-I model as an empirical improvement for completeness. All Chandra spectral fits reported in this paper were performed in the 0.7 to 7.0 keV range.
@@
$\alpha 1.490 Al K$$\alpha$
$\alpha 1.786 Si K$$\alpha$
2.141 Au $\alpha_{1}/M$$\alpha_{2}$ 2.142 Au $\alpha_{1}/M$$\alpha_{2}$
2.6∗∗ - - -
$\alpha 7.478 Ni K$$\alpha$
@@
9.74 Au $\alpha_{1}/L$$\alpha_{2}$ 9.740 Au $\alpha_{1}/L$$\alpha_{2}$
- - 9.755 Au $\alpha_{1}/L$$\alpha_{2}$